OpenStudy (sepeario):

The average of one group of numbers is 4. A second group contains twice as many numbers and has an average of 10. What is the average of both groups of numbers?

OpenStudy (mokeira):

\[\frac{ 10+4 }{ 2 }\] im just guessing

OpenStudy (sepeario):

Sorry @Mokeira that is wrong the answer is 8.

OpenStudy (mokeira):

could you explain to me how

OpenStudy (sepeario):

That's what I'm asking lol! @Mokeira

OpenStudy (mokeira):

lol...lemme ask a couple of people

OpenStudy (mokeira):

@aryandecoolest

OpenStudy (mokeira):

@midhun.madhu1987

OpenStudy (mokeira):

@hopelovelift

OpenStudy (mokeira):

@ganeshie8

OpenStudy (mokeira):

@Preetha

OpenStudy (mokeira):

@thomaster

OpenStudy (anonymous):

I would love to help however it seems like I am not capable? Is there anymore data available? e,g what numbers are in group one and two. It feels like the questions is lacking values.

OpenStudy (sepeario):

Yeah, @hopelovelift thats it unfortunately.

OpenStudy (mokeira):

let the total of numbers in the group be x so the get the total multiply average by x =4x The other group has a total of 2x members so total when you multiply by 10 is 20x and total members is x+2x=3x so new average is \[\frac{ (20x+4x) }{ 3x }\] Ans=8 hope you have understood

OpenStudy (anonymous):

LOOL nvm, I was soo close as well. You bet me too it mokeira

OpenStudy (mokeira):

hahaha...yaay!!! That was a great question!

OpenStudy (anonymous):

Not going to lie, I had to manually input numbers in first. For example 4+4/2=4 10+20+5+5/4=10 Then I realized theres a pattern... This question took me a while. Was fun solving though.

OpenStudy (anonymous):

Oh just now saw the question...anyways you guys are brilliant :)